OpenStudy (nudeoua):

Given the function g(x) = −3x + 4, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. The value of g(−2) is smaller than the value of g(4). The value of g(−2) is the same as the value of g(4). The values of g(−2) and g(4) cannot be compared. The value of g(−2) is larger than the value of g(4).

OpenStudy (anonymous):

Don't be confused by the different variable, if it helps replace the g with an f. Then just plug the given number into the equation for the value of x and solve. ie g(-2) = -3(-2) + 4 and g(4) = -3(4) + 4 g(-2) = ? and g(4) = ? Compare those two answers to the criteria given

OpenStudy (nudeoua):

g(-2)= 10 g(4)= 8 ?

OpenStudy (anonymous):

-3 x 4 = -12 -> add 4 to -12 and you get -8

OpenStudy (anonymous):

10 is right for the first one, good job.